Kine, fancy footwear manufacturing company has an obligation to pay MXN 14 million in 30 days for a recent shipment from Mexico. The CFO of Kine is contemplating hedging the company’s MXN exposure on this transaction. She collects the below information regarding the interest rates and exchange rates, from her forex trader: Spot Rate: MXN 20.08 / USD Forward Rate: MXN 20.28 / USD 30-day Put Option on USD MXN 19.50 / USD: 1% Premium 30-day Call Option on USD MXN 20.50/ USD: 3% Premium USD 30-day interest rate (annualized): 7.5% MXN 30-day interest rate (annualized): 15%


1
Expert's answer
2021-11-15T17:12:02-0500

A) SCL can use Forward, Call option, and money market.

Forward cost:

"14000000/20.08—14000000/20.28=6875.85\ USD"

Options cost:

"14000000/20.08-14000000/19.5-14000000/19.5*0.01=-27917\ USD"

b) "14000000/(1+0.15*30/365)/20.08=688719\ USD"

"688719*(1+0.075*30/365)=692964\ USD"

"14000000/20.08-692964=4247\ USD"